OpenStudy (anonymous):

Maclaurin Series e^(x/2)

OpenStudy (anonymous):

Use the classical series for e^x, but replace x with x/2. That's all. Straightforward.

OpenStudy (anonymous):

we got this assignment but we dont have topic for any series

OpenStudy (anonymous):

e^x = 1 + x + x^2/2! + x^3/3! + x^4/4! + .... so, e^(x/2) = 1+ (x/2) + (x/2)^2/2! + (x/2)^3/3! + ....

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

Welcome